Calculate the slope of a line, that passes through the origin, and the mid-point of the segment joining the points P (0, -4) and B (8, 0).

Grade:12

1 Answers

SJ
3 years ago
Welcome to Askiitians

Given that,

The coordinates of the mid-point of the line segment joining the points P (0, -4) and B (8, 0) are:[(0+8)/2 , (-4+0)/2] = (4, -2)It is known that the slope (m) of a non-vertical line passing through the points (x1, y1) and (x2,

y2) is given by the formula

m = (y2-y1) / ( (x2-x1), where (x2is not equal to x1)

Therefore, the slope of the line passing through the points (0, 0,) and (4, -2) is

m= (-2-0)/(4-0)

m= -2/4

m= -½

Hence, the required slope of the line is -1/2

Regards